# Implicit Bias of Mirror Descent for Shallow Neural Networks in Univariate Regression

### v1

@@ A.1 Linearization
Consider network (1) with ReLU or absolute value activation function and initialization (2). Assume that the potential function satisfies Assumption 1, and that the training data $\{x_{i}\}_{i\in[m]}$ satisfies $x_{i}\neq x_{j}$ when $i\neq j$ and $\{x_{i}\}_{i=1}^{m}\subset\operatorname{supp}(p_{\mathcal{B}})$. Then there exists a constant $\lambda_{0}>0$, which is determined by the activation function, the training data, and the initialization distribution, such that $\operatorname{plim}_{n\rightarrow\infty}\lambda_{\min}(H(0))=\operatorname{plim}_{n\rightarrow\infty}\lambda_{\min}(\tilde{H}(0))=(\phi^{\prime\prime}(0))^{-1}\lambda_{0}$.
@@
$=\frac{1}{n\phi^{\prime\prime}(0)}\sum_{k=1}^{n}\sigma(w_{k}(0)x_{i}-b_{k}(0))\sigma(w_{k}(0)x_{j}-b_{k}(0))+\frac{1}{n\phi^{\prime\prime}(0)}$
$\quad\quad+\frac{1}{n^{2}\phi^{\prime\prime}(0)}\sum_{k=1}^{n}(\sqrt{n}a_{k}(0))^{2}\sigma^{\prime}(w_{k}(0)x_{i}-b_{k}(0))\sigma^{\prime}(w_{k}(0)x_{j}-b_{k}(0))(x_{i}x_{j}+1).$
Note $\{\sqrt{n}a_{k}\}_{k\in[n]}$ can be viewed as independent samples from the random variable $\mathcal{A}$. Therefore, by law of large number, we have
$\operatorname{plim}_{n\rightarrow\infty}H_{ij}(0)=\frac{1}{\phi^{\prime\prime}(0)}\mathbb{E}\left[\sigma(\mathcal{W}x_{i}-\mathcal{B})\sigma(\mathcal{W}x_{j}-\mathcal{B})\right].$
Define $H^{\infty}=(\phi^{\prime\prime}(0))^{-1}G$ where $G_{ij}=\mathbb{E}[\sigma(\mathcal{W}x_{i}-\mathcal{B})\sigma(\mathcal{W}x_{j}-\mathcal{B})]$. Then we have $\operatorname{plim}_{n\rightarrow\infty}H(0)=H^{\infty}$. In the meantime, we have
@@
with inner-product $\langle h,h^{\prime}\rangle_{\mathcal{H}}=\mathbb{E}[h(\mathcal{W},\mathcal{B})h^{\prime}(\mathcal{W},\mathcal{B})]$. For each $x_{i}\in\mathbb{R}$, define $h_{i}(w,b)=\sigma(wx_{i}-b)$. Since $(\mathcal{W},\mathcal{B})$ is sub-Gaussian and $h_{i}$ is uniformly continuous, $h_{i}\in\mathcal{H}$. Note $G$ is the Gram matrix of $\{h_{i}\}_{i=1}^{m}$. Hence, $G\succ 0$ is equivalent to $\{h_{i}\}_{i=1}^{m}$ are linearly independent in $\mathcal{H}$.
Assume $\sum_{i=1}^{m}c_{i}h_{i}=0$ in $\mathcal{H}$, which means $\sum_{i=1}^{m}c_{i}h_{i}(w,b)=0$ almost everywhere. Note $\{h_{i}\}_{i=1}^{m}$ are continuous. It follows that $\sum_{i=1}^{m}c_{i}h_{i}(w,b)=0$ everywhere. Consider a function $t\colon\operatorname{supp}(p_{\mathcal{B}})\rightarrow\mathbb{R}$, defined by $t(b)=\sum_{i=1}^{m}c_{i}h_{i}(1,b)$. Now fix an $i\in[m]$. Since $t(b)\equiv 0$, $t(b)$ is differentiable at $b=x_{i}$. Also notice $h_{j}(1,b)=\sigma(x_{j}-b)$ is differentiable at $b=x_{i}$ if and only if $j\neq i$. Therefore, we must have $c_{i}=0$. Noticing $i$ is arbitrary, we have the claimed results. ∎
In the next proposition, we show that the model output $\mathbf{f}(t)$ converges exponentially fast to $\mathbf{y}$ and the parameters have $O_{p}(n^{-1/2})$ bounded $l_{\infty}$ distance from the initialzation.
@@
Consider a single-hidden-layer univariate network (1) with ReLU or absolute value activation function and initialization (2). Assume the number of hidden neurons is sufficiently large. Given any finite training data set $\{(x_{i},y_{i})\}_{i=1}^{m}$ that satisfies $x_{i}\neq x_{j}$ when $i\neq j$ and $\{x_{i}\}_{i=1}^{m}\subset\operatorname{supp}(p_{\mathcal{B}})$, consider mirror flow (4) with the potential satisfying Assumptions 1 and learning rate $\eta=\eta_{0}/n$ with $\eta_{0}>0$. Then there exist constants $C_{1},C_{2}>0$ such that with high probability over initialization the following holds:
@@
$\|\mathbf{f}(t)-\mathbf{y}\|_{2}\leq\exp(-\eta_{0}C_{2}t)\|\mathbf{f}(0)-\mathbf{y}\|_{2},\ \forall t\geq 0.$
@@
$C_{1}=\frac{8\phi^{\prime\prime}(0)\beta\sqrt{m}(B_{x}+1)K\|\mathbf{f}(0)-\mathbf{y}\|_{2}}{\lambda_{0}},\quad C_{2}=\frac{\lambda_{0}}{4\phi^{\prime\prime}(0)},$ (20)
where $\lambda_{0}=\operatorname{plim}_{n\rightarrow\infty}\lambda_{\min}(\frac{1}{n}J_{\theta}\mathbf{f}(0)J_{\theta}\mathbf{f}(0)^{T})$, $\beta=\max_{z\in[-1,1]}(\phi^{\prime\prime}(z))^{-1}$, $m$ is the size of the training data set, $B_{x}=\max_{i\in[m]}|x_{i}|$, and $K$ is a constant determined by $\mathcal{B}$ and $\mathcal{A}$.
@@
[Proof of Proposition 13] To ease the notation, we let $J(t)=J_{\theta}\mathbf{f}(t)$ and $Q(t)=(\nabla^{2}\Phi(\theta(t)))^{-1}$. For $i\in[m]$ and $k\in[n]$, let $\sigma_{ik}(t)=\sigma(w_{k}(t)x_{i}-b_{k}(t))$ and $\sigma^{\prime}_{ik}(t)=\sigma^{\prime}(w_{k}(t)x_{i}-b_{k}(t))$.
@@
$\tau(t)=\max\Big{\{}\sqrt{n}\|\boldsymbol{a}(t)-\boldsymbol{a}(0)\|_{\infty},n\|\boldsymbol{w}(t)-\boldsymbol{w}(0)\|_{\infty},n\|\boldsymbol{b}(t)-\boldsymbol{b}(0)\|_{\infty},n|d(t)-d(0)|\Big{\}},$
@@
In the following let $t$ denote any point in $[0,t_{1}]$. Noticing $\tau(t)\leq C_{1}$, we have for large enough $n$,
$\|Q(t)\|_{2}=\|\operatorname{Diag}(\frac{1}{\phi^{\prime\prime}(\theta_{k}(t)-\theta_{k}(0))})\|_{2}=\max_{k\in[n]}\frac{1}{\phi^{\prime\prime}(\theta_{k}(t)-\theta_{k}(0))}\leq\max_{z\in[-1,1]}\frac{1}{\phi^{\prime\prime}(z)}=\beta.$
@@
$\begin{dcases}I_{1}=\frac{1}{n}\sum_{k=1}^{n}\Big{|}\frac{\sigma_{ik}(t)\sigma_{jk}(t)}{\phi^{\prime\prime}(a_{k}(t)-a_{k}(0))}-\frac{\sigma_{ik}(0)\sigma_{jk}(0)}{\phi^{\prime\prime}(0)}\Big{|};\\ I_{2}=|x_{i}x_{j}|\frac{1}{n}\sum_{k=1}^{n}\Big{|}\frac{a_{k}(t)^{2}\sigma^{\prime}_{ik}(t)\sigma^{\prime}_{jk}(t)}{\phi^{\prime\prime}(w_{k}(t)-w_{k}(0))}-\frac{a_{k}(0)^{2}\sigma^{\prime}_{ik}(0)\sigma^{\prime}_{jk}(0)}{\phi^{\prime\prime}(0)}\Big{|};\\ I_{3}=\frac{1}{n}\sum_{k=1}^{n}\Big{|}\frac{a_{k}(t)^{2}\sigma^{\prime}_{ik}(t)\sigma^{\prime}_{jk}(t)}{\phi^{\prime\prime}(b_{k}(t)-b_{k}(0))}-\frac{a_{k}(0)^{2}\sigma^{\prime}_{ik}(0)\sigma^{\prime}_{jk}(0)}{\phi^{\prime\prime}(0)}\Big{|};\\ I_{4}=\frac{1}{n}\Big{|}\frac{1}{\phi^{\prime\prime}(d(t)-d(0))}-\frac{1}{\phi^{\prime\prime}(0)}\Big{|}.\end{dcases}$ (22)
Notice that according to Assumption (1), $\frac{1}{\phi^{\prime\prime}(\cdot)}$ is continuous and hence is uniformly continuous on $[-1,1]$. Then noticing $\|\theta(t)-\theta(0)\|_{\infty}=O(n^{-1/2})$, we have for large enough $n$,
$\max_{k\in[n]}\Big{|}\frac{1}{\phi^{\prime\prime}(\theta_{k}(t)-\theta_{k}(0))}-\frac{1}{\phi^{\prime\prime}(0)}\Big{|}=o_{p}(1).$ (23)
@@
$\operatorname{plim}_{n\rightarrow\infty}\frac{1}{n}\sum_{k\in[n]}|\sigma_{ik}\sigma_{jk}|=E[\sigma(\mathcal{W}x_{i}-\mathcal{B})\sigma(\mathcal{W}x_{j}-\mathcal{B})].$
@@
$I_{2}$ $\leq\frac{B_{x}^{2}}{n}\sum_{k\in[n]}\Big{(}a_{k}(0)^{2}\beta|\sigma^{\prime}_{ik}(t)\sigma^{\prime}_{jk}(t)-\sigma^{\prime}_{ik}(0)\sigma^{\prime}_{ik}(0)|$
$\quad+a_{k}(0)^{2}\Big{|}\frac{1}{\phi^{\prime\prime}(w_{k}(t)-w_{k}(0))}-\frac{1}{\phi^{\prime\prime}(0)}\Big{|}+|a_{k}(t)^{2}-a_{k}(0)^{2}|\beta\Big{)}$
$=\frac{B_{x}^{2}}{n}\Big{(}2\beta\|\boldsymbol{a}(0)\|_{2}^{2}+o_{p}(1)\|\boldsymbol{a}(0)\|_{2}^{2}+\beta\sum_{k\in[n]}o_{p}(1)\Big{)}$
@@
$\frac{\mathrm{d}}{\mathrm{dt}}\|\mathbf{f}(t)-\mathbf{y}\|_{2}^{2}\leq-\frac{\eta_{0}\lambda_{0}}{2\phi^{\prime\prime}(0)}\|\mathbf{f}(t)-\mathbf{y}\|_{2}^{2}$
@@
$\|\mathbf{f}(t)-\mathbf{y}\|_{2}\leq\exp(-\frac{\eta_{0}\lambda_{0}}{4\phi^{\prime\prime}(0)}t)\|\mathbf{f}(0)-y\|_{2}.$ (26)
@@
$\|J_{a_{k}}\mathbf{f}(t)\|_{2}\leq\sqrt{m}\|J_{a_{k}}\mathbf{f}\|_{\infty}\leq\sqrt{m}(B_{x}|w_{k}(t)|+|b_{k}(t)|)\leq 2K(B_{x}+1)\sqrt{mn}.$ (27)
@@
$|\frac{\mathrm{d}}{\mathrm{dt}}a_{k}(t)|$ $\leq\frac{\eta_{0}}{n}\|Q(t)\|_{2}\|J_{a_{k}}\mathbf{f}(t)\|_{2}\|\mathbf{f}(t)-\mathbf{y}\|_{2}$ (28)
$\leq\frac{\eta_{0}}{n}\beta\cdot 2K(B_{x}+1)\sqrt{mn}\cdot\exp(-\frac{\eta_{0}\lambda_{0}}{4\phi^{\prime\prime}(0)}t)\|\mathbf{f}(0)-\mathbf{y}\|_{2}.$
@@
$|a_{k}(t)-a_{k}(0)|\leq\int_{0}^{t}|\frac{\mathrm{d}}{\mathrm{dt}}a_{k}(s)|\mathrm{d}s<\frac{8\phi^{\prime\prime}(0)\beta\sqrt{m}(B_{x}+1)K\|\mathbf{f}(0)-\mathbf{y}\|_{2}}{\lambda_{0}}n^{-1/2}=C_{1}n^{-1/2}.$ (29)
@@
$\begin{dcases}\|J_{w_{k}}\mathbf{f}(t)\|_{2}\leq\sqrt{m}B_{x}|a_{k}(t)|\leq 2K\sqrt{m}B_{x},\quad\forall k\in[n];\\ \|J_{b_{k}}\mathbf{f}(t)\|_{2}\leq\sqrt{m}|a_{k}(t)|\leq 2K\sqrt{m},\quad\forall k\in[n];\\ \|J_{d}\mathbf{f}(t)\|_{2}=\sqrt{m}.\end{dcases}$ (30)
@@
$\|\tilde{\mathbf{f}}(t)-\mathbf{y}\|_{2}\leq\exp(-\eta_{0}C_{2}t)\|\mathbf{f}(0)-\mathbf{y}\|_{2},\ \forall t\geq 0.$
@@
Under the same assumption in Proposition 13, given any $x\in\mathbb{R}$, with high probability over initialization,
@@
[Proof of Proposition 15] Under initialization 2, we have $\|\boldsymbol{a}(0)\|_{2}=O_{p}(1)$ and thus
$|H_{ij}(0)-\tilde{H}_{ij}(0)|$ $\leq\frac{1}{n}\sum_{k\in[n]}|a_{k}(0)|^{2}|\sigma^{\prime}_{ik}(0)\sigma^{\prime}_{jk}(0)||x_{i}x_{j}+1|+\frac{1}{n\phi^{\prime\prime}(0)}$ (31)
$\leq\frac{B_{x}^{2}+1}{n}\|\boldsymbol{a}(0)\|_{2}+O_{p}(n^{-1})$

### v2

@@ A.1 Linearization
Consider network (1) with ReLU or absolute value activation function and initialization (2). Assume that the potential function satisfies Assumption 1, and that the training data $\{x_{i}\}_{i\in[m]}$ satisfies $x_{i}\neq x_{j}$ when $i\neq j$, and $\{\|x_{i}\|_{2}\}_{i=1}^{m}\subset\operatorname{supp}(p_{\mathcal{B}})$. Then there exists a constant $\lambda_{0}>0$, which is determined by the activation function, the training data, and the initialization distribution, such that $\operatorname{plim}_{n\rightarrow\infty}\lambda_{\min}(H(0))=\operatorname{plim}_{n\rightarrow\infty}\lambda_{\min}(\tilde{H}(0))=(\phi^{\prime\prime}(0))^{-1}\lambda_{0}$.
@@
$=\frac{1}{n\phi^{\prime\prime}(0)}\sum_{k=1}^{n}\sigma(\hat{w}_{k}^{T}x_{i}-\hat{b}_{k})\sigma(\hat{w}_{k}^{T}x_{j}-\hat{b}_{k})+\frac{1}{n\phi^{\prime\prime}(0)}$
$\quad\quad+\frac{1}{n\phi^{\prime\prime}(0)}\sum_{k=1}^{n}\hat{a}_{k}^{2}\sigma^{\prime}(\hat{w}_{k}^{T}x_{i}-\hat{b}_{k})\sigma^{\prime}(\hat{w}_{k}^{T}x_{j}-\hat{b}_{k})(x_{i}^{T}x_{j}+1).$
Noticing $\{a_{k}\}_{k\in[n]}$ are independent samples from the random variable $\frac{1}{\sqrt{n}}\mathcal{A}$ and by sub-Gaussian concentration bound, we have that $\max_{k\in[n]}|\hat{a}_{k}|\leq C\sqrt{(\log n)/n}$ holds with high probability for some constant $C>0$. It then follows that:
$\operatorname{plim}_{n\rightarrow\infty}H_{ij}(0)=\frac{1}{\phi^{\prime\prime}(0)}\mathbb{E}\left[\sigma(\mathcal{W}^{T}x_{i}-\mathcal{B})\sigma(\mathcal{W}^{T}x_{j}-\mathcal{B})\right].$
Define $H^{\infty}=(\phi^{\prime\prime}(0))^{-1}G$ where $G_{ij}=\mathbb{E}[\sigma(\mathcal{W}^{T}x_{i}-\mathcal{B})\sigma(\mathcal{W}^{T}x_{j}-\mathcal{B})]$. Then we have $\operatorname{plim}_{n\rightarrow\infty}H(0)=H^{\infty}$. In the meantime, we have
@@
with inner-product $\langle h,h^{\prime}\rangle_{\mathcal{H}}=\mathbb{E}[h(\mathcal{W},\mathcal{B})h^{\prime}(\mathcal{W},\mathcal{B})]$. For each $x_{i}\in\mathbb{R}$, define $h_{i}(w,b)=\sigma(w^{T}x_{i}-b)$. Since $(\mathcal{W},\mathcal{B})$ is sub-Gaussian and $h_{i}$ is uniformly continuous, $h_{i}\in\mathcal{H}$. Note $G$ is the Gram matrix of $\{h_{i}\}_{i=1}^{m}$. Hence, $G\succ 0$ is equivalent to $\{h_{i}\}_{i=1}^{m}$ are linearly independent in $\mathcal{H}$.
Assume $\sum_{i=1}^{m}c_{i}h_{i}=0$ in $\mathcal{H}$, which means $\sum_{i=1}^{m}c_{i}h_{i}(w,b)=0$ almost everywhere. Without loss of generality, assume $\|x_{1}\|_{2}\geq\|x_{2}\|_{2}\geq\ldots\geq\|x_{m}\|\geq 0$. We aim to show $c_{1}=0$. Notice that, for any $i\in[m]$, $h_{i}$ is not differentiable at $(w,b)$ if and only if $w^{T}x_{i}-b=0$. In particular, $h_{1}$ is not differentiable at $(x_{1}/\|x_{1}\|_{2},\|x_{1}\|_{2})$. Now for all $j>1$, by noticing that $\|x_{j}\|\leq\|x_{1}\|$ and $x_{1}\neq x_{j}$ we have,
$(\frac{x_{1}}{\|x_{1}\|_{2}})^{T}x_{j}-\|x_{1}\|_{2}=\|x_{j}\|\cos(\angle(x_{1},x_{j}))-\|x_{1}\|_{2}<0$
where $\angle(x_{1},x_{j})$ denotes the angle between $x_{1}$ and $x_{j}$. This implies that $h_{j}$ is differentiable at $(x_{1}/\|x_{1}\|_{2},\|x_{1}\|_{2})$ and so is $\sum_{j\neq 1}c_{j}h_{j}$. Therefore, to make $\sum_{i}c_{i}h_{i}$ differentiable everywhere, we have $c_{1}=0$. With the same strategy, we can show $c_{j}=0$ for $j=2,3,\ldots,m-1$, since $\|x_{j}\|>0$. It then follows that all coefficients are zero, which completes the proof. ∎
In the next proposition, we show that the model output $\boldsymbol{f}(t)$ converges exponentially fast to $\boldsymbol{y}$ and the parameters have $O_{p}(n^{-1/2})$ bounded $\ell_{\infty}$ distance from the initialzation.
@@
Consider a single-hidden-layer univariate network (1) with ReLU or absolute value activation function and initialization (2). Assume the number of hidden neurons is sufficiently large. Given any finite training data set $\{(x_{i},y_{i})\}_{i=1}^{m}$ that satisfies $x_{i}\neq x_{j}$ when $i\neq j$, and $\{\|x_{i}\|_{2}\}_{i=1}^{m}\subset\operatorname{supp}(p_{\mathcal{B}})$, consider mirror flow (4) with the potential satisfying Assumptions 1 and learning rate $\eta=\eta_{0}/n$ with $\eta_{0}>0$. Then there exist constants $C_{1},C_{2}>0$ such that with high probability over initialization the following holds:
@@
$\|\boldsymbol{f}(t)-\boldsymbol{y}\|_{2}\leq\exp(-\eta_{0}C_{2}t)\|\boldsymbol{f}(0)-\boldsymbol{y}\|_{2},\ \forall t\geq 0.$
@@
$C_{1}=\frac{8\phi^{\prime\prime}(0)\beta\sqrt{m}(B_{x}+1)K\|\boldsymbol{f}(0)-\boldsymbol{y}\|_{2}}{\lambda_{0}},\quad C_{2}=\frac{\lambda_{0}}{4\phi^{\prime\prime}(0)},$ (20)
where $\lambda_{0}=\operatorname{plim}_{n\rightarrow\infty}\lambda_{\min}(\frac{1}{n}J_{\theta}\boldsymbol{f}(0)J_{\theta}\boldsymbol{f}(0)^{T})$, $\beta=\max_{z\in[-1,1]}(\phi^{\prime\prime}(z))^{-1}$, $m$ is the size of the training data set, $B_{x}=\max_{i\in[m]}\|x_{i}\|_{2}$, and $K$ is a constant determined by $\mathcal{B}$ and $\mathcal{A}$.
@@
[Proof of Proposition 14] To ease the notation, we let $J(t)=J_{\theta}\boldsymbol{f}(t)$ and $Q(t)=(\nabla^{2}\Phi(\theta(t)))^{-1}$. For $i\in[m]$ and $k\in[n]$, let $\sigma_{ik}(t)=\sigma(w_{k}(t)^{T}x_{i}-b_{k}(t))$ and $\sigma^{\prime}_{ik}(t)=\sigma^{\prime}(w_{k}(t)^{T}x_{i}-b_{k}(t))$.
@@
$\tau(t)=\max\Big{\{}\sqrt{n}\|\boldsymbol{a}(t)-\boldsymbol{a}(0)\|_{\infty},n\max_{k\in[n]}\|w_{k}(t)-w_{k}(0)\|_{\infty},n\|\boldsymbol{b}(t)-\boldsymbol{b}(0)\|_{\infty},n|d(t)-d(0)|\Big{\}},$
@@
In the following let $t$ denote any point in $[0,t_{1}]$. Noticing that $\tau(t)\leq C_{1}$ implies $\|\theta(t)-\theta(0)\|_{\infty}\leq C_{1}n^{-1/2}$, we have for large enough $n$,
$\|Q(t)\|_{2}=\|\operatorname{Diag}(\frac{1}{\phi^{\prime\prime}(\theta_{k}(t)-\theta_{k}(0))})\|_{2}=\max_{k\in[3n+1]}\frac{1}{\phi^{\prime\prime}(\theta_{k}(t)-\theta_{k}(0))}\leq\max_{z\in[-1,1]}\frac{1}{\phi^{\prime\prime}(z)}=\beta.$ (21)
@@
$\begin{dcases}I_{1}=\frac{1}{n}\sum_{k=1}^{n}\Big{|}\frac{\sigma_{ik}(t)\sigma_{jk}(t)}{\phi^{\prime\prime}(a_{k}(t)-a_{k}(0))}-\frac{\sigma_{ik}(0)\sigma_{jk}(0)}{\phi^{\prime\prime}(0)}\Big{|};\\ I_{2}=\frac{1}{n}\sum_{k=1}^{n}\sum_{r=1}^{d}|x_{i,r}x_{j,r}|\Big{|}\frac{a_{k}(t)^{2}\sigma^{\prime}_{ik}(t)\sigma^{\prime}_{jk}(t)}{\phi^{\prime\prime}(w_{k,r}(t)-w_{k,r}(0))}-\frac{a_{k}(0)^{2}\sigma^{\prime}_{ik}(0)\sigma^{\prime}_{jk}(0)}{\phi^{\prime\prime}(0)}\Big{|};\\ I_{3}=\frac{1}{n}\sum_{k=1}^{n}\Big{|}\frac{a_{k}(t)^{2}\sigma^{\prime}_{ik}(t)\sigma^{\prime}_{jk}(t)}{\phi^{\prime\prime}(b_{k}(t)-b_{k}(0))}-\frac{a_{k}(0)^{2}\sigma^{\prime}_{ik}(0)\sigma^{\prime}_{jk}(0)}{\phi^{\prime\prime}(0)}\Big{|};\\ I_{4}=\frac{1}{n}\Big{|}\frac{1}{\phi^{\prime\prime}(d(t)-d(0))}-\frac{1}{\phi^{\prime\prime}(0)}\Big{|}.\end{dcases}$ (23)
Notice that according to Assumption (1), $\frac{1}{\phi^{\prime\prime}(\cdot)}$ is continuous and hence is uniformly continuous on $[-1,1]$. Then by noticing $\|\theta(t)-\theta(0)\|_{\infty}=O(n^{-1/2})$, we have for large enough $n$,
$\max_{k\in[3n+1]}\Big{|}\frac{1}{\phi^{\prime\prime}(\theta_{k}(t)-\theta_{k}(0))}-\frac{1}{\phi^{\prime\prime}(0)}\Big{|}=o_{p}(1).$ (24)
@@
$\operatorname{plim}_{n\rightarrow\infty}\frac{1}{n}\sum_{k\in[n]}|\sigma_{ik}\sigma_{jk}|=E[\sigma(\mathcal{W}^{T}x_{i}-\mathcal{B})\sigma(\mathcal{W}^{T}x_{j}-\mathcal{B})].$
@@
$I_{2}$ $\leq\frac{B_{x}^{2}}{n}\sum_{k,r}\Big{(}a_{k}(0)^{2}\beta|\sigma^{\prime}_{ik}(t)\sigma^{\prime}_{jk}(t)-\sigma^{\prime}_{ik}(0)\sigma^{\prime}_{ik}(0)|$
$\quad+a_{k}(0)^{2}\Big{|}\frac{1}{\phi^{\prime\prime}(w_{k,r}(t)-w_{k,r}(0))}-\frac{1}{\phi^{\prime\prime}(0)}\Big{|}+|a_{k}(t)^{2}-a_{k}(0)^{2}|\beta\Big{)}$
$=\frac{B_{x}^{2}d}{n}\Big{(}2\beta\|\boldsymbol{a}(0)\|_{2}^{2}+o_{p}(1)\|\boldsymbol{a}(0)\|_{2}^{2}+\beta n\cdot o_{p}(1)\Big{)}$
@@
$\frac{\mathrm{d}}{\mathrm{dt}}\|\boldsymbol{f}(t)-\boldsymbol{y}\|_{2}^{2}\leq-\frac{\eta_{0}\lambda_{0}}{2\phi^{\prime\prime}(0)}\|\boldsymbol{f}(t)-\boldsymbol{y}\|_{2}^{2}$
@@
$\|\boldsymbol{f}(t)-\boldsymbol{y}\|_{2}\leq\exp(-\frac{\eta_{0}\lambda_{0}}{4\phi^{\prime\prime}(0)}t)\|\boldsymbol{f}(0)-y\|_{2}.$ (27)
@@
$\|J_{a_{k}}\boldsymbol{f}(t)\|_{2}\leq\sqrt{m}\|J_{a_{k}}\boldsymbol{f}\|_{\infty}\leq\sqrt{m}(B_{x}\|w_{k}(t)\|_{2}+|b_{k}(t)|)\leq 2K(B_{x}+1)\sqrt{mn}.$ (28)
@@
$|\frac{\mathrm{d}}{\mathrm{dt}}a_{k}(t)|$ $\leq\frac{\eta_{0}}{n}\|Q(t)\|_{2}\|J_{a_{k}}\boldsymbol{f}(t)\|_{2}\|\boldsymbol{f}(t)-\boldsymbol{y}\|_{2}$ (29)
$\leq\frac{\eta_{0}}{n}\beta\cdot 2K(B_{x}+1)\sqrt{mn}\cdot\exp(-\frac{\eta_{0}\lambda_{0}}{4\phi^{\prime\prime}(0)}t)\|\boldsymbol{f}(0)-\boldsymbol{y}\|_{2}.$
@@
$|a_{k}(t)-a_{k}(0)|\leq\int_{0}^{t}|\frac{\mathrm{d}}{\mathrm{dt}}a_{k}(s)|\mathrm{d}s<\frac{8\phi^{\prime\prime}(0)\beta\sqrt{m}(B_{x}+1)K\|\boldsymbol{f}(0)-\boldsymbol{y}\|_{2}}{\lambda_{0}}n^{-1/2}=C_{1}n^{-1/2}.$ (30)
@@
$\begin{dcases}\|J_{w_{k,r}}\boldsymbol{f}(t)\|_{2}\leq\sqrt{m}B_{x}|a_{k}(t)|\leq 2K\sqrt{m}B_{x},\quad\forall k\in[n],r\in[d];\\ \|J_{b_{k}}\boldsymbol{f}(t)\|_{2}\leq\sqrt{m}|a_{k}(t)|\leq 2K\sqrt{m},\quad\forall k\in[n];\\ \|J_{d}\boldsymbol{f}(t)\|_{2}=\sqrt{m}.\end{dcases}$ (31)
@@
$\|\tilde{\boldsymbol{f}}(t)-\boldsymbol{y}\|_{2}\leq\exp(-\eta_{0}C_{2}t)\|\boldsymbol{f}(0)-\boldsymbol{y}\|_{2},\ \forall t\geq 0.$
@@
Under the same assumption in Proposition 14, given any $x\in\mathbb{R}^{d}$, with high probability over initialization,
@@
[Proof of Proposition 16] For $i\in[m]$ and $k\in[n]$, let $\sigma^{\prime}_{ik}(t)=\sigma^{\prime}(w_{k}(t)^{T}x_{i}-b_{k}(t))$. Under initialization 2, we have $\|\boldsymbol{a}(0)\|_{2}=O_{p}(1)$ and thus
$|H_{ij}(0)-\tilde{H}_{ij}(0)|$ $\leq\frac{1}{n\phi^{\prime\prime}(0)}\sum_{k\in[n]}|a_{k}(0)|^{2}|\sigma^{\prime}_{ik}(0)\sigma^{\prime}_{jk}(0)||x_{i}^{T}x_{j}+1|+\frac{1}{n\phi^{\prime\prime}(0)}$ (32)
$\leq\frac{B_{x}^{2}+1}{n\phi^{\prime\prime}(0)}\|\boldsymbol{a}(0)\|_{2}+O_{p}(n^{-1})$